June 11th, 2015 (Evening Run)

Aim: Calibrating TOFs, testing Run Control, testing Online Reconstruction

Daily schedule

June 11th
Time 9pm - 5am
MOM V. Blackmore
Lead shifter S. Middleton
Shifter 2 J. Nugent
BLOC A. Dobbs
SOC A. Dobbs
Experimenter A. Dobbs, P. Hanlet
Physics Shifter P. Franchini
Beam? Yes
Additional Personnel

Shift plan:

  • 7:30pm
    • Shifters + BLOC arrive
    • Hall and beam line to be closed and prepared for beam
  • 9pm
    • E. Overton to remove tracker veto from DAQ before TOF calibration begins
    • Check TOF1 response channel-by-channel
    • Operate experiment via Run Control, take data through to TOF2/KL/EMR to calibrate TOFs and test Online Reconstruction
    • Settings as A. Dobbs requires

Data Requested

  • TOF calibration

Beamline Configurations

  • We expect to run with 2V beamloss and the beam bump active.
  • DS is currently off
  • Beam settings via Run Control
    • TOF calibration, short runs for Run Control testing
    • TOF calibration, long 100k triggers run -- aim for 1000 triggers/pixel (~25'000 spills, or until the end of the shift)

TOF Calibration run

Number of pulses Beamloss Particle Species p at Tgt p@D1 p@D2 Proton Absorber Q1 Q2 Q3 D1 DS D2 Q4 Q5 Q6 Q7 Q8 Q9 Similar Run(s)
# V mm A A A A A A A A A A A A A A A A DAQ run #
25000 min 1 Positron 300.0 259.7 249.58 29 56.35 102.7 62.65 195.72 --- 98.59 163.16 218.79 114.14 83.40 125.70 80.06 1678

Phone details for 6pm daily meeting

Phone access : ReadyTalk number

US 866-740-1260

UK 0800-496-0576

Swiss 0800-70-53-51

Italy 800182592

Access Code: 2155402

Updated by Blackmore, Victoria about 9 years ago ยท 6 revisions